THE ROTATIONAL SPECTRUM OF H^{15}NO_{3}: ALL STATES BELOW 1000 cm^{−1}

\underline{DOUGLAS T. PETKIE}, MARK KIPLING, and ASHELY JONES, Department of Physics, Wright State University, Dayton OH 45435; PAUL HELMINGER, Department of Physics, University of South Alabama, Mobile, AL 36688; IVAN MEDVEDEV and ATSUKO MAEDA, Department of Physics, The Ohio State University, Columbus, OH 43210; BRIAN J. DROUIN and CHARLES E. MILLER, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

The rotational spectrum of H^{15}NO_{3} was recorded using isotope enriched samples at Ohio State University with the FASSST spectrometer and at the Jet Propulsion Laboratory with the cascaded frequency multiplication spectrometer. The OSU system used a heated cell over the frequency range of 118-370 GHz while the JPL room temperature measurements included the frequency ranges of 74-109, 400-410, 639-656, and 800-850 GHz. Transitions in the ground and six lowest vibrational states, 6^{1}, 7^{1}, 8^{1}, 9^{1}, and the 5^{1}/9^{2} dyad, have been assigned and fit using Watson-type Hamiltonians. The 9^{1} and 9^{2} states require torsional parameters to account for the observed torsional splitting of ∼2.4 MHz and ∼70 MHz, respectively. Fermi and Coriolis interactions were included to accurately describe the strong interactions in the 5^{1}/9^{2} dyad and to account for an observed torsional splitting of ∼15 MHz induced onto the 5^{1} state. The analysis of each state will be presented along with a discussion of the spectroscopic constants.